# KVQuant: Towards 10 Million Context Length LLM Inference with KV Cache Quantization

### v4

@@ Abstract.
LLMs are seeing growing use for applications which require large context windows, and with these large context windows KV cache activations surface as the dominant contributor to memory consumption during inference. Quantization is a promising approach for compressing KV cache activations; however, existing solutions fail to represent activations accurately in sub-4-bit precision. Our work, KVQuant, facilitates low precision KV cache quantization by incorporating several novel methods: (i) Per-Channel Key Quantization, where we adjust the dimension along which we quantize the Key activations to better match the distribution; (ii) Pre-RoPE Key Quantization, where we quantize Key activations before the rotary positional embedding to mitigate its impact on quantization; (iii) Non-Uniform KV Cache Quantization, where we derive per-layer sensitivity-weighted non-uniform datatypes that better represent the distributions; and (iv) Per-Vector Dense-and-Sparse Quantization, where we isolate outliers separately for each vector to minimize skews in quantization ranges. By applying our method to the LLaMA, Llama-2, Llama-3, and Mistral models, we achieve $<0.1$ perplexity degradation with 3-bit quantization on both Wikitext-2 and C4, outperforming existing approaches. Our method enables serving LLaMA-7B with a context length of up to 1 million on a single A100-80GB GPU and up to 10 million on an 8-GPU system. We develop custom CUDA kernels for KVQuant, showing that we can achieve up to $\sim$1.7$\times$ speedups, compared to baseline fp16 matrix-vector multiplications, for the LLaMA-7B model. The code is available at https://github.com/SqueezeAILab/KVQuant/.
1. Introduction
@@
Large language models (LLMs) have revolutionized many natural language processing (NLP) tasks. In order to improve the capabilities of LLMs, there is significant interest in increasing the context lengths of LLMs. Longer context lengths enable new applications, including long document summarization, retrieval for answering questions about long documents, extended multi-turn applications (chen2023longlora, ), and code analysis. To support this pull from applications, there have been significant recent advances in long-context length models in industry (anthropic, ; openai, ), as well as in academia (chen2023longlora, ; liu2024world, ).
@@
Given the importance of LLM workloads, there is strong motivation to improve their inference efficiency. LLM inference with large context lengths can be incredibly resource-intensive; serving LLMs requires high-end GPUs, and the largest LLMs require costly multi-GPU inference setups. When analyzing the computational nature of generative inference with LLMs, it becomes quickly apparent that, for relatively small batch sizes, the computation is memory bound (kim2023squeezellm, ). With the growing divergence between computational speeds and memory speeds, this problem is only going to get worse over time (gholami2020ai_and_memory_wall, ). This makes reducing the memory bottleneck preeminently important. Further analysis shows that the memory bottleneck is strongly related to context size. For short sequence lengths, the dominant contributor to memory consumption is the weight matrices, and therefore the optimal strategy is to minimize the model size in order to reduce memory consumption as well as bandwidth requirements (kim2023full, ; kim2023squeezellm, ). However, as shown in Figure 1, the main bottleneck for long sequence lengths is the memory requirements for caching Key and Value (KV) activations throughout inference. This challenge is further exacerbated when one considers batched inference.
@@
Figure 1. Overview of the different components used in KVQuant that result in less than 0.1 perplexity degradation over the fp16 baseline when quantizing the KV cache for the LLaMA-7B model to 3-bit precision. As shown in  Table 2, our 3-bit approach results in 4.8×4.8$\times$4.8 × reduction in cached activation memory footprint.
Table 1. Model size and activation memory size for different sequence lengths and batch sizes (BS) for different LLaMA models. For long sequence lengths and larger batch sizes, activation memory is the main bottleneck (particularly when weights are already quantized to low precision). At a sequence length of 128K with the LLaMA-7B model, the KV cache is the main bottleneck (see below left). Additionally, if model weights are quantized, then the KV cache is the main bottleneck even at a sequence length of 32K. By compressing the KV cache to 2-bit precision, we can enable 1M context length inference with the LLaMA-7B model on a single A100-80GB GPU, and we can also enable 10M context length inference with the LLaMA-7B model on an 8-GPU system.
BS Model Model Size (GB) fp16 KV Cache Size with Different Seq. Len. (GB)
16 $\rightarrow$ 2-bit 32K 128K 1M 10M (16 $\rightarrow$ 2-bit)
1 7B 12.6 $\rightarrow$ 1.6 16 64 512 4883 $\rightarrow$ 610
13B 24.1 $\rightarrow$ 3.0 25 100 800 7629 $\rightarrow$ 954
30B 60.3 $\rightarrow$ 7.5 49 195 1560 14877 $\rightarrow$ 1860
65B 121.1 $\rightarrow$ 15.1 80 320 2560 24414 $\rightarrow$ 3052
4 7B 12.6 $\rightarrow$ 1.6 64 256 2048 19531 $\rightarrow$ 2441
13B 24.1 $\rightarrow$ 3.0 100 400 3200 30518 $\rightarrow$ 3815
30B 60.3 $\rightarrow$ 7.5 195 780 6240 59509 $\rightarrow$ 7439
65B 121.1 $\rightarrow$ 15.1 320 1280 10240 97656 $\rightarrow$ 12207
2. Background
2.1. LLM Inference
When inferring a decoder-only LLM, inference proceeds in two distinct phases. In the prefill phase, the model takes in an input prompt, which it processes in parallel. During the generation phase, the model then generates the output sequence autoregressively, meaning that each token generation is dependent on all previously generated tokens. As such, for small batch sizes, the generation phase of LLM inference is typically memory-bandwidth bound, as the only available parallelism is across different sequences in a given batch.
@@ 2.1. LLM Inference
Additionally, during generation, the model needs to store intermediate Key and Value activations in order to condition generations on previously generated output tokens. Otherwise, we would need to recompute all prior Keys and Values at each timestep, which would be prohibitively expensive. For each prior token, we need to store the Keys and Values at each layer in order to use these activations when generating future tokens. These stored activations are referred to as the Key-Value (KV) cache. Throughout this paper, we will capitalize Key and Value to distinguish when we are referring to the KV cache tensors. Assuming a model with $n$ layers and $h$ attention heads with dimension $d$ that is stored using $e$ bytes per element, the KV cache size for batch size $b$ and sequence length $l$ is $2\cdot n\cdot h\cdot d\cdot e\cdot b\cdot l$, meaning that it grows linearly with both batch size and sequence length. As shown in Table 1, the KV cache becomes the dominant contributor to memory consumption for longer sequence lengths and larger batch sizes. Note that since each sequence in batched inference depends on separate past context, there is no available batch-level parallelism when loading the cached Keys and Values for their respective computations in batched inference. KV cache loading is therefore always memory-bandwidth bound. This motivates pursuing methods to optimally compress the KV cache, even at the expense of a more complex dequantization process.
2.2. LLM Quantization
@@
KV Cache Quantization. There have been many prior works on LLM quantization. Several have focused on weight-only quantization for LLMs, due to the greater contribution to memory consumption and runtime for fairly small sequence length and batch size (lin2023awq, ; dettmers2023spqr, ; kim2023squeezellm, ). There has also been work on quantizing both weights and activations (including KV cache) (xiao2023smoothquant, ; shao2023omniquant, ). However, there is still a significant perplexity degradation when quantizing KV cache activations to low precision; (sheng2023flexgen, ; zhao2023atom, ) quantized KV cache activations to 4-bits, but required fine-grained grouping for 4-bit quantization, while still observing some perplexity degradation, and (sheng2023flexgen, ) observed that 3-bit KV cache quantization with fine-grained grouping leads to unacceptable accuracy loss. Other works quantized KV cache activations to 4-bits but required retraining to maintain performance (liu2023llmqat, ). One concurrent work also explores low precision KV cache quantization in order to enable larger batch size inference by reducing the KV cache size (kivi, ). In this work, we introduce a method for near-lossless low-bit KV cache quantization that minimizes performance degradation without the need for finetuning.
Outlier-Aware LLM Quantization. LLMs have been known to have distinct outliers both in weights and activations (dettmers2022llm, ; dettmers2023spqr, ; kim2023squeezellm, ). SqueezeLLM and SpQR both decompose the weight matrix into a sparse matrix containing a small portion of outliers and a dense matrix that can be accurately quantized to low precision (referred to as dense-and-sparse or sparse-quantized representation) (dettmers2023spqr, ; kim2023squeezellm, ). LLM.int8() (dettmers2022llm, ) handled particular outlier channels separately in higher precision, and SmoothQuant (xiao2023smoothquant, ) migrates quantization difficulty due to outlier channels to weights in order to support joint weight-activation quantization. Other works reconsidered the dimension along which we quantize in order to reduce quantization error (or else added per-channel compensation to improve quantization performance) (bondarenko2021understanding, ; heo2023rethinking, ; wei2022outlier, ; wei2023outlier, ). In this work, we demonstrate that per-channel pre-RoPE Key quantization provides significant accuracy benefits given the outlier structure in Keys, and that dense-and-sparse quantization can be efficiently applied for KV cache quantization.
Non-uniform LLM Quantization. Non-uniform quantization has also been applied in the context of LLMs. Non-uniform quantization allows for more flexible quantization signpost placement relative to uniform quantization methods, enabling improved accuracy for the same bit precision (kim2023squeezellm, ; dettmers2023qlora, ). Building on the observation that model parameters tend to be approximately normally-distributed, prior work has proposed the NormalFloat datatype (dettmers2023qlora, ). SqueezeLLM (kim2023squeezellm, ) derived per-channel non-uniform quantization signposts using a sensitivity-weighted k-means approach. In this work, we show that we can derive accurate per-layer non-uniform datatypes using a sensitivity-weighted k-means approach with KV cache activations.
2.3. KV Cache Compression
@@
There have also been several prior works on compressing the KV cache. Some of these methods aim to only store important tokens in the KV cache and to evict less important tokens, thereby maintaining low memory usage (liu2023scissorhands, ; zhang2023h, ; ge2023model, ; li2024snapkv, ). Other methods aim to only retrieve a subset of tokens at each step to achieve memory bandwidth savings (ribar2023sparq, ). In this work, we explore KV cache quantization as an orthogonal direction for compressing the KV cache in order to enable long context inference.
@@
Figure 2. Example distributions of the activation values for Keys pre-RoPE, Keys post-RoPE, and Values for LLaMA-7B on a sample with 2K sequence length from Wikitext-2. We observe several patterns: (i) Keys pre-RoPE exhibit clear outliers in specific channels across different tokens; (ii) after applying RoPE, the distribution becomes less structured and there are less consistent magnitudes for outlier channels (this is expected, as RoPE applies a rotation operation between pairs of channels); and (iii) Values exhibit no fixed outlier pattern with outlier values across channels and tokens.
3. Method
3.1. Per-Channel Key Quantization
@@
To inform our approach, we first performed a detailed analysis to understand the KV cache distributions. Figure 2 shows sample distributions for the KV cache activations. We observe that the Key matrices tend to have distinct outlier channels, which have larger average magnitudes than other channels; this corroborates previous observations about outlier channels in LLM activations (dettmers2022llm, ; xiao2023smoothquant, ). The Value matrices exhibit both outlier channels as well as outlier tokens (although these outliers are less extreme than the outlier Key channels).
@@
Existing KV cache quantization approaches perform per-token quantization (meaning that the scaling factor and zero-point are shared by elements in the same token) (sheng2023flexgen, ; zhao2023atom, ). However, due to the differing average magnitudes between channels, the values within a channel are easier to quantize when grouped together than the values across different channels. As such, to better match the distributions, we investigate per-channel KV cache quantization, meaning that the scaling factor and zero-point are shared by elements in the same channel. By sharing the scaling factor and zero-point along the channel dimension, this will naturally group together values with similar magnitudes, thereby mitigating the impacts of outlier channels on other channels when quantizing to low precision. As outlined in Appendix E, we find that per-channel quantization provides significant accuracy benefits for Keys but not for Values. By leveraging per-channel quantization for Keys and per-token quantization for Values, we observe a 3.82 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization. Note that this can potentially add runtime overhead since the quantization dimension is now misaligned with the reduction dimension for the Keys when performing matrix-vector multiplications. However, we find that we are able to efficiently dequantize Keys and perform the Query-Key matrix-vector multiplication without adding runtime overhead, as shown in Section 4.4. Additionally, as outlined in Section 3.6, per-channel quantization can also be challenging due to the need to recompute scaling factors as tokens are added to the Key cache. We show that we can calibrate offline for scaling factors, thereby avoiding expensive online recomputation.
@@
Per-channel Key quantization was also explored in another concurrent work (kivi, ), which leveraged similar intuition about grouping together large magnitude values in the same channel to minimize quantization error. Their methodology requires fine-grained grouping for per-channel quantization while maintaining a residual subset of the KV cache in fp16. In our work, we demonstrate that by leveraging offline calibration, we can accurately perform per-channel quantization without grouping.
3.2. Pre-RoPE Key Quantization
One issue when quantizing Keys is handling the rotary positional embedding (RoPE), which is applied to Keys and Queries in most public LLMs, including LLaMA and Llama-2 (su2023roformer, ). Given Query and Key vectors $Q_{m}=W_{q}*x_{m}$ and $K_{n}=W_{k}*x_{n}$ at positions $m$ and $n$ in the sequence, RoPE is applied as position-dependent rotations to each of these vectors to obtain $\tilde{Q}_{m}$ and $\tilde{K}_{n}$. This embeds the relative position between a Query and Key vector as an amount of an angle that is a multiple of its position index. Formally, RoPE is applied in self-attention as follows:
(1) $\tilde{Q}_{m}\tilde{K}_{n}^{\top}$ $=(R^{d}_{\theta,m}\cdot Q_{m})(R^{d}_{\theta,n}\cdot K_{n})^{\top}.$
@@ 3.2. Pre-RoPE Key Quantization
When caching Key vectors, we therefore need to either cache $\tilde{K}_{n}$, or else we need to cache $K_{n}$ and apply $R^{d}_{\theta,n}$ on-the-fly during inference. The challenge with caching Key vectors after applying this rotation is that it leads to mixing pairs of channels by different amounts for different positions in the sequence, as shown in Appendix A (since it jointly rotates pairs of channels by different angles depending on the position in the sequence). The post-RoPE activation distribution is also shown in Figure 2, demonstrating how the rotation between pairs of channels leads to less consistent channel magnitudes. This makes it harder to quantize Key activation channels which would typically have consistent large-magnitude values. This motivated our investigation into whether we could perform pre-RoPE Key quantization (meaning that we quantize $K_{n}$) and then efficiently apply the positional embeddings on-the-fly after dequantization. The benefits of pre-RoPE Key quantization are highlighted in Appendix F, yielding 0.82 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization. To be able to quantize Keys pre-RoPE, we develop a fused kernel to efficiently apply RoPE post-dequantization (the details of this approach will be discussed in Section 3.7).
3.3. nuqX: An X-Bit Per-Layer Sensitivity-Weighted Non-Uniform Datatype
Figure 3. One typically achieves better performance when the scaling factor/zero point are computed online. However, this is quite challenging to do for per-channel quantization, as these factors will not only need to be recomputed for every new Key appended to the Key cache, but also all the prior cached Keys will need to be updated. As such, we use a calibration set to compute per-channel scaling factors offline. A similar challenge exists for per-token quantization, but online calibration for this does not require updating prior cached Values. In Section 3.6 and Appendix J, we discuss how we are able to efficiently compute outlier thresholds / scaling factors for per-token calibration, thereby enabling online computation.
@@
Uniform quantization is suboptimal for KV cache quantization since the Query and Key activations are non-uniform. Additionally, KV cache loading is memory bandwidth bound, regardless of batch size or sequence length, meaning that the dequantization overhead introduced by non-uniform quantization methods is not problematic (since the added computation does not introduce any additional latency). It is therefore desirable to leverage non-uniform quantization methods for KV cache quantization. In (kim2023squeezellm, ), the authors computed non-uniform quantization signposts using a sensitivity-weighted k-means approach. However, this cannot be applied directly to KV cache quantization as the Values are quantized dynamically at runtime, which means that we would need to apply K-means online during inference, and it is also difficult to estimate sensitivity for activations online.
We therefore facilitate efficient online non-uniform KV cache quantization by deriving a per-tensor non-uniform datatype offline on calibration data, which is then rescaled per-channel or per-token to accurately represent the key and value distributions. We compute sensitivity-weighted quantization signposts offline on a calibration set prior to inference, while maintaining compatibility with per-vector quantization by separately normalizing each channel to the range $[-1,1]$ prior to deriving the shared datatype. Using the diagonal Fisher information matrix (derived in Appendix B), along with the quantization error for activation $A$, we formulate the error minimization objective as follows, where $A$ is flattened to one dimension and where $N$ is the number of elements from all of the samples in our calibration set:
(2) $Q(A)^{*}\simeq\operatorname*{arg\,min}_{Q}\sum_{i=1}^{N}\mathcal{F}_{ii}\big{(}A_{i}-Q(A_{i})\big{)}^{2}.$
@@
We modify the objective in Equation 2 as described in Appendix C in order to apply it using the normalized activation values. We then minimize it offline on a calibration set using a k-means solver in order to obtain the quantization signposts for the non-uniform datatype for each Key or Value layer. Appendix G compares our non-uniform quantization approach with existing uniform and non-uniform quantization baselines (dettmers2023qlora, ), demonstrating how our non-uniform approach provides 0.29 perplexity improvement on Wikitext-2 for LLaMA-7B relative to 3-bit uniform methods. Table 14 in Appendix J shows how computing the required Fisher information for the LLaMA-65B model takes only a few minutes, and how using the k-means solver takes only a few minutes per layer (with the computation for each layer being parallelizable).
3.4. Per-Vector Dense-and-Sparse Quantization

### v5

@@ Abstract
LLMs are seeing growing use for applications which require large context windows, and with these large context windows KV cache activations surface as the dominant contributor to memory consumption during inference. Quantization is a promising approach for compressing KV cache activations; however, existing solutions fail to represent activations accurately in sub-4-bit precision. Our work, KVQuant, facilitates low precision KV cache quantization by incorporating several novel methods: (i) Per-Channel Key Quantization, where we adjust the dimension along which we quantize the Key activations to better match the distribution; (ii) Pre-RoPE Key Quantization, where we quantize Key activations before the rotary positional embedding to mitigate its impact on quantization; (iii) Non-Uniform KV Cache Quantization, where we derive per-layer sensitivity-weighted non-uniform datatypes that better represent the distributions; and (iv) Per-Vector Dense-and-Sparse Quantization, where we isolate outliers separately for each vector to minimize skews in quantization ranges. By applying our method to the LLaMA, Llama-2, Llama-3, and Mistral models, we achieve $<0.1$ perplexity degradation with 3-bit quantization on both Wikitext-2 and C4, outperforming existing approaches. Our method enables serving LLaMA-7B with a context length of up to 1 million on a single A100-80GB GPU and up to 10 million on an 8-GPU system. We develop custom CUDA kernels for KVQuant, showing that we can achieve up to $\sim$1.7$\times$ speedups, compared to baseline fp16 matrix-vector multiplications, for the LLaMA-7B model. Code is available at https://github.com/SqueezeAILab/KVQuant.
1 Introduction
@@
Large language models (LLMs) have revolutionized many natural language processing (NLP) tasks. In order to improve the capabilities of LLMs, there is significant interest in increasing the context lengths of LLMs. Longer context lengths enable new applications, including long document summarization, retrieval for answering questions about long documents, extended multi-turn applications [6], and code analysis. To support this pull from applications, there have been significant recent advances in long-context length models in industry [2, 30], as well as in academia [6, 22].
@@
Given the importance of LLM workloads, there is strong motivation to improve their inference efficiency. LLM inference with large context lengths can be incredibly resource-intensive; serving LLMs requires high-end GPUs, and the largest LLMs require costly multi-GPU inference setups. When analyzing the computational nature of generative inference with LLMs, it becomes quickly apparent that, for relatively small batch sizes, the computation is memory bound [17]. With the growing divergence between computational speeds and memory speeds, this problem is only going to get worse over time [12]. This makes reducing the memory bottleneck preeminently important. Further analysis shows that the memory bottleneck is strongly related to context size. For short sequence lengths, the dominant contributor to memory consumption is the weight matrices, and therefore the optimal strategy is to minimize the model size in order to reduce memory consumption as well as bandwidth requirements [18, 17]. However, as shown in Figure 1, the main bottleneck for long sequence lengths is the memory requirements for caching Key and Value (KV) activations throughout inference. This challenge is further exacerbated when one considers batched inference.
@@
Figure 1: Left: Model size versus activation memory size for the LLaMA-7B model with sequence length 512 and 128K. For longer context lengths, the KV cache becomes the dominant memory bottleneck. Memory consumption of model weights and KV cache activations for different LLaMA models with different sequence lengths are provided in Table 7 in Appendix A. Right: Overview of the different components used in KVQuant that result in less than 0.1 perplexity degradation over the fp16 baseline when quantizing the KV cache for the LLaMA-7B model to 3-bit precision. As shown in  Table 1, our 3-bit approach results in 4.8×4.8$\times$4.8 × reduction in cached activation memory footprint.
2 Background
2.1 LLM Inference
@@
When inferring a decoder-only LLM, inference proceeds in two distinct phases. In the prefill phase, the model takes in an input prompt, which it processes in parallel. During the generation phase, the model then generates the output sequence autoregressively, meaning that each token generation is dependent on all previously generated tokens. As such, for small batch sizes, the generation phase of LLM inference is typically memory-bandwidth bound, as the only available parallelism is across different sequences in a given batch. Additionally, during generation, the model needs to store intermediate Key and Value activations at each layer in order to condition generations on previously generated output tokens. Otherwise, we would need to recompute all prior Keys and Values at each timestep, which would be prohibitively expensive. These stored activations are referred to as the Key-Value (KV) cache. Throughout this paper, we will capitalize Key and Value to distinguish when we are referring to the KV cache tensors. Assuming a model with $n$ layers and $h$ attention heads with dimension $d$ that is stored using $e$ bytes per element, the KV cache size for batch size $b$ and sequence length $l$ is $2\cdot n\cdot h\cdot d\cdot e\cdot b\cdot l$, meaning that it grows linearly with both batch size and sequence length. As shown in Table 7, the KV cache becomes the dominant contributor to memory consumption for longer sequence lengths and larger batch sizes. Note that since each sequence in batched inference depends on separate past context, there is no available batch-level parallelism when loading the cached Keys and Values for their respective computations in batched inference. KV cache loading is therefore always memory-bandwidth bound. This motivates pursuing methods to optimally compress the KV cache, even at the expense of a more complex dequantization process.
2.2 LLM Quantization
There have been many prior works on LLM quantization. Several have focused on weight-only quantization for LLMs, due to the greater contribution to memory consumption and runtime for small sequence lengths and batch sizes [21, 9, 17]. Prior works have noted the presence of distinct outliers in both weights and activations [7, 9, 17]. One approach that has been developed to address this outlier issue in the context of weight quantization is dense-and-sparse quantization, where each weight matrix is decomposed into a sparse outlier matrix and a dense low-precision matrix [9, 17]. Prior works have also leveraged non-uniform quantization methods to improve accuracy for the same bit precision by allowing for flexible quantization signpost placement [17, 8]. These approaches have either used a fixed non-uniform datatype such as NormalFloat [8], or derived quantization signposts using a sensitivity-weighted k-means approach [17]. Appendix B provides a more detailed overview of related work for outlier-aware LLM quantization and non-uniform LLM quantization.
@@ 2.2 LLM Quantization
There has also been work on quantizing both weights and activations (including KV cache) [41, 33]. However, there is still a significant perplexity degradation when quantizing KV cache activations to low precision; [34, 44] quantized KV cache activations to 4-bits, but required fine-grained grouping for 4-bit quantization, while still observing some perplexity degradation, and [34] observed that 3-bit KV cache quantization with fine-grained grouping leads to unacceptable accuracy loss. Other works quantized KV cache activations to 4-bits but required retraining to maintain performance [24]. One concurrent work also explores low precision KV cache quantization in order to enable larger batch size inference by reducing the KV cache size [26].
2.3 KV Cache Compression
@@
There have also been several prior works on compressing the KV cache. Some of these methods aim to only store important tokens in the KV cache and to evict less important tokens, thereby maintaining low memory usage [25, 43, 11, 20]. Other methods aim to only retrieve a subset of tokens at each step to achieve memory bandwidth savings [32]. In this work, we explore KV cache quantization as an orthogonal direction for compressing the KV cache in order to enable long context inference.
@@
Figure 2: Example distributions of the activation values for Keys pre-RoPE, Keys post-RoPE, and Values for LLaMA-7B on a sample with 2K sequence length from Wikitext-2. We observe several patterns: (i) Keys pre-RoPE exhibit clear outliers in specific channels across different tokens; (ii) after applying RoPE, the distribution becomes less structured and there are less consistent magnitudes for outlier channels (this is expected, as RoPE applies a rotation operation between pairs of channels); and (iii) Values exhibit no fixed outlier pattern with outlier values across channels and tokens.
3 Method
3.1 Per-Channel Key Quantization
@@
To inform our approach, we first performed a detailed analysis to understand the KV cache distributions. Figure 2 shows sample distributions for the KV cache activations. We observe that the Key matrices tend to have distinct outlier channels, which have larger average magnitudes than other channels; this corroborates previous observations about outlier channels in LLM activations [7, 41]. The Value matrices exhibit both outlier channels as well as outlier tokens (although these outliers are less extreme than the outlier Key channels).
@@
Existing KV cache quantization approaches perform per-token quantization (meaning that the scaling factor and zero-point are shared by elements in the same token) [34, 44]. However, due to the differing average magnitudes between channels, the values within a channel are easier to quantize when grouped together than the values across different channels. As such, to better match the distributions, we investigate per-channel KV cache quantization, meaning that the scaling factor and zero-point are shared by elements in the same channel. By sharing the scaling factor and zero-point along the channel dimension, this will naturally group together values with similar magnitudes, thereby mitigating the impacts of outlier channels on other channels when quantizing to low precision. As outlined in Appendix G, we find that per-channel quantization provides significant accuracy benefits for Keys but not for Values. By leveraging per-channel quantization for Keys and per-token quantization for Values, we observe a 3.82 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization. Note that this can potentially add runtime overhead since the quantization dimension is now misaligned with the reduction dimension for the Keys when performing matrix-vector multiplications. However, we find that we are able to efficiently dequantize Keys and perform the Query-Key matrix-vector multiplication without adding runtime overhead, as shown in Section 4.4. Additionally, as outlined in Section 3.6, per-channel quantization can also be challenging due to the need to recompute scaling factors as tokens are added to the Key cache. We show that we can calibrate offline for scaling factors, thereby avoiding expensive online recomputation.
@@
Per-channel Key quantization was also explored in another concurrent work [26], which leveraged similar intuition about grouping together large magnitude values in the same channel to minimize quantization error. Their methodology requires fine-grained grouping for per-channel quantization while maintaining a residual subset of the KV cache in fp16. In our work, we demonstrate that by leveraging offline calibration, we can accurately perform per-channel quantization without grouping.
3.2 Pre-RoPE Key Quantization
@@
One issue when quantizing Keys is handling the rotary positional embedding (RoPE), which is applied to Keys and Queries in most public LLMs, including LLaMA and Llama-2 [35]. Given Query and Key vectors $Q_{m}=W_{q}*x_{m}$ and $K_{n}=W_{k}*x_{n}$ at positions $m$ and $n$ in the sequence, RoPE is applied as position-dependent rotations to each of these vectors to obtain $\tilde{Q}_{m}=R^{d}_{\theta,m}\cdot Q_{m}$ and $\tilde{K}_{n}=R^{d}_{\theta,n}\cdot K_{n}$. This embeds the relative position between a Query and Key vector as an amount of an angle that is a multiple of its position index. When caching Key vectors, we therefore need to either cache $\tilde{K}_{n}$, or else we need to cache $K_{n}$ and apply $R^{d}_{\theta,n}$ on-the-fly during inference. The challenge with caching Key vectors after applying this rotation is that it leads to mixing pairs of channels by different amounts for different positions in the sequence, as shown in Appendix C (since it jointly rotates pairs of channels by different angles depending on the position in the sequence). The post-RoPE activation distribution is also shown in Figure 2, demonstrating how the rotation between pairs of channels leads to less consistent channel magnitudes. This makes it harder to quantize Key activation channels which would typically have consistent large-magnitude values. This motivated our investigation into whether we could perform pre-RoPE Key quantization (meaning that we quantize $K_{n}$) and then efficiently apply the positional embeddings on-the-fly after dequantization. The benefits of pre-RoPE Key quantization are highlighted in Appendix H, yielding 0.82 perplexity improvement on Wikitext-2 for 3-bit LLaMA-7B quantization. To be able to quantize Keys pre-RoPE, we develop a fused kernel to efficiently apply RoPE post-dequantization (the details of this approach will be discussed in Section 3.7).
3.3 nuqX: An X-Bit Per-Layer Sensitivity-Weighted Non-Uniform Datatype
@@
Uniform quantization is suboptimal for KV cache quantization since the Query and Key activations are non-uniform. Additionally, KV cache loading is memory bandwidth bound, regardless of batch size or sequence length, meaning that the dequantization overhead introduced by non-uniform quantization methods is not problematic (since the added computation does not introduce any additional latency). It is therefore desirable to leverage non-uniform quantization methods for KV cache quantization. In [17], the authors computed non-uniform quantization signposts using a sensitivity-weighted k-means approach. However, this cannot be applied directly to KV cache quantization as the Values are quantized dynamically at runtime, which means that we would need to apply K-means online during inference, and it is also difficult to estimate sensitivity for activations online. We therefore facilitate efficient online non-uniform KV cache quantization by deriving a per-tensor non-uniform datatype offline on calibration data, which is then rescaled per-channel or per-token to accurately represent the key and value distributions. We compute sensitivity-weighted quantization signposts offline on a calibration set prior to inference, while maintaining compatibility with per-vector quantization by separately normalizing each channel to the range $[-1,1]$ prior to deriving the shared datatype. Using the diagonal Fisher information matrix (derived in Appendix D), along with the quantization error for activation $A$, we formulate the error minimization objective as follows, where $A$ is flattened to one dimension and where $N$ is the number of elements from all of the samples in our calibration set:
$Q(A)^{*}\simeq\operatorname*{arg\,min}_{Q}\sum_{i=1}^{N}\mathcal{F}_{ii}\big{(}A_{i}-Q(A_{i})\big{)}^{2}.$ (1)
@@
We modify the objective in Equation 1 as described in Appendix E in order to apply it using the normalized activation values. We then minimize it offline on a calibration set using a k-means solver in order to obtain the quantization signposts for the non-uniform datatype for each Key or Value layer. Appendix I compares our non-uniform quantization approach with existing uniform and non-uniform quantization baselines [8], demonstrating how our non-uniform approach provides 0.29 perplexity improvement on Wikitext-2 for LLaMA-7B relative to 3-bit uniform methods. Table 16 in Appendix L shows how computing the required Fisher information for the LLaMA-65B model takes only a few minutes, and how using the k-means solver takes only a few minutes per layer (with the computation for each layer being parallelizable).
3.4 Per-Vector Dense-and-Sparse Quantization